# Imprints of an early matter-dominated era arising from dark matter dilution mechanism on cosmic string dynamics and gravitational wave signatures

### v2

@@ I Introduction
Modern cosmology originates from the study of the universe expansion and element formation [1]. Further observations give rise to the standard $\Lambda\mathrm{CDM}$ cosmological model, which suggests that after inflation and reheating, the universe successively experiences a radiation-dominated (RD) era, a matter-dominated era (MD), and a dark-energy-dominated era. Main evidences involve the observations of the cosmic expansion [2], the big bang nucleosynthesis (BBN) [3, 4], the cosmic microwave background [5, 6, 7, 8], and the accelerating expansion at late times [9, 10].
Nevertheless, these observational results can hardly date back to pre-BBN eras. Therefore, it is essential to maintain an open perspective regarding the early history of the universe before BBN [11]. Various hypotheses beyond the standard cosmic history have been put forward, such as an early matter-dominated (EMD) era [12, 13, 14, 15, 16, 17, 18, 19], a kination-dominated era [20, 21, 22, 23, 24, 25], and an intermediate inflationary era [26, 27, 28, 29]. Verification of such novel hypotheses about the cosmic history necessitates observational probes capable of accessing epochs prior to BBN, when the universe was opaque to photons and traditional electromagnetic detection methods are ineffective. Nonetheless, this challenge can be addressed through observations of gravitational waves (GWs), a new messenger first detected in 2015 [30]. Unlike electromagnetic radiation, GWs can propagate freely through space, preserving information from the early universe and reaching us in the present day.
@@
In this work, we explore how the SGWB spectrum originated from a preexisting CS network is modified by an EMD era that arises within the framework of the dark matter (DM) dilution mechanism [59, 60, 61, 62, 63, 64, 65, 66, 67, 68]. This mechanism addresses the DM overproduction problem by relying on significant entropy production from the decays of a long-lived particle, referred to as the “dilutor”. For the dilution mechanism to be effective, the dilutor must dominate the energy density of the universe for a finite period, thereby introducing an EMD era that interrupts the conventional RD era. To quantify the impact of this EMD era, we will trace the evolution of the total energy density of the universe by solving the relevant Boltzmann equations. This allows us to derive the resulting SGWB spectrum, which reflects the influence of the EMD era on the cosmic string dynamics and the GW emissions.
This paper is outlined as follows. In Section II, we review the formation of CS loops and analyze their number densities during RD and MD eras, which are essential for assessing the SGWB spectrum induced by CS loops. In Section III, we investigate an EMD era emerging from the DM dilution mechanism and examine its effects on the evolution of the CS network, the loop number densities, and the resulting SGWB spectrum. Finally, in Section IV, we summarize our findings.
@@ II Gravitational waves originating from cosmic strings
In a scalar field theory with a global or gauge $\mathrm{U}(1)$ symmetry, CSs could be formed after the spontaneous breaking of the $\mathrm{U}(1)$ symmetry in the early universe. They are one-dimensional topological defects concentrating energy of the scalar field (and the gauge field for the gauged case) [69]. In the Nambu-Goto approximation, CSs are described as infinitely thin objects with tension $\mu$, which is the energy per unit length. The dimensionless quantity $G\mu$ is commonly used to represent the CS tension. Considering the dynamics of CSs, these combined factors establish CSs as promising sources of GWs [70, 35].
@@ II.1 Gravitational waves from cosmic string loops
Cosmic strings are generated randomly in the early universe, leading to the formation of a CS network. Long strings with super-horizon lengths could intersect with each other to form CS loops, whose relativistic oscillations can effectively emit GWs. Although small structures in long strings can also produce GWs, their contributions are generally negligible, compared to those generated by loops [71, 72, 37]. In practice, it is sufficient to consider only the stochastic GWs originating from CS loops.
@@
where the coefficient $\Gamma$ is estimated to be about 50 [74], and $G$ represents the Newtonian gravitational constant. The frequencies of GWs emitted by a CS loop of length $l$ are [70]
@@
where $P_{n}$ is the dimensionless emission power in units of $G\mu^{2}$ for a single mode $n$, which can be estimated by numerical simulations for GW emissions in RD and MD eras [74].
@@
Note that this result corresponds to the GW emission time $t_{\mathrm{e}}$, but GW experiments only receive signals at the present time $t_{0}$. Hence, the effect of the cosmological redshift must be accounted for, and the present GW frequency is given by $f=a(t_{\mathrm{e}})f_{\mathrm{e}}$, while the present GW energy density is $\rho_{\mathrm{GW}}=\rho_{\mathrm{GW}}(t_{\mathrm{e}})a^{4}(t_{\mathrm{e}})$, where $a(t)$ is the scale factor normalized such that $a(t_{0})=1$. Integrating Eq. (5) over time, we arrive at
@@
is a function of the redshift $z$ introduced to account for the changes in relativistic degrees of freedom, where $g_{\star}(z)$ and $g_{\star s}(z)$ are the effective numbers of relativistic degrees of freedom for the energy and entropy densities, respectively. Considering the evolution of $g_{\star}(z)$ and $g_{\star s}(z)$, we can approximate $\mathcal{G}(z)$ as a piecewise function which changes its value at the epochs of electron-positron annihilation $(z\simeq 10^{9})$ and QCD phase transition $(z\simeq 2\times 10^{12})$[75]
@@
However, if an EMD era exists, the evolution of the Hubble rate $H(t)$ would be altered, necessitating modifications to the expressions above. This situation will be discussed in detail in Section III.
@@ II.2 Number density of cosmic string loops
The evolution of the CS network can be appropriately described using the velocity-dependent one-scale (VOS) model [72, 77, 78], where the correlation length $L$ and the root-mean-square (RMS) velocity $v$ of string segments serve as two fundamental parameters characterizing the dynamics of the network. Thus, the energy density of long strings can be expressed as $\rho=\mu/L^{2}$, and its evolution in the universe is governed by [77]
@@
where the friction effect is neglected. The term $-2H\rho$ accounts for the dilution and stretching of the strings due to the cosmic expansion, while the term $-2Hv^{2}\rho$ captures the energy loss caused by the redshifting of velocities. The term $-\tilde{c}v\rho/L$ represents the energy removed through loop formation, where $\tilde{c}\simeq 0.23$ [79] is the loop chopping efficiency. Additionally, the evolution of the RMS velocity $v$ is determined by [77]
@@
Introducing a dimensionless quantity $\xi\equiv L/t$, which represents the correlation length $L$ normalized by the cosmic time $t$, the evolution equations above can be rewritten as
@@
By numerically solving these equations, one finds that the solutions rapidly converge to constant values of $\xi$ and $v$, independent of the initial conditions. This indicates that the CS network quickly evolves into a linear scaling regime characterized by $L\propto t$. By setting $\dot{\xi}=\dot{v}=0$, the scaling solutions for RD and MD eras can be derived as [80]
@@
where the Lorentz factor $\gamma_{v}\equiv(1-v^{2})^{-1/2}$ incorporates the energy loss caused by the redshifting of the loop velocity, and $\mathcal{F}$ is a coefficient that characterizes the fraction of slow, large loops responsible for the dominant contribution to GW emissions. The values of $\mathcal{F}$ differ between RD and MD eras, denoted as $\mathcal{F}_{\mathrm{r}}$ and $\mathcal{F}_{\mathrm{m}}$, respectively. These values can be determined by comparison with numerical simulations in the scaling regime [82, 83].
@@
quantifies the rate of increase in the number density of CS loops per unit length per unit time. Because of GW emission, a CS loop of length $l^{\prime}$ at time $t^{\prime}$ loses energy $\Gamma G\mu^{2}(t-t^{\prime})$ by time $t$ and consequently shrinks to a length $l=l^{\prime}-\Gamma G\mu(t-t^{\prime})$. Therefore, the loop number density per unit loop length at time $t$ is given by
@@
with $l^{\prime}(t)=l+\Gamma G\mu(t-t^{\prime})$.$t_{\mathrm{ini}}$ denotes the initial time of loop production, and the scale factors account for the dilution because of cosmic expansion. To determine $n_{\mathrm{CS}}(l,t)$ for calculating the energy density of emitted GWs according to Eq. (4), knowledge of the loop production function $\mathcal{P}(l,t)$ is required.
During an RD era, numerical simulations [83] reveal that most large loops are generated with a length $l$ characterized by a fraction $\alpha_{\mathrm{r}}=l/L$ of the correlation length $L$. $\alpha_{\mathrm{r}}$ satisfies $\alpha_{\mathrm{r}}\xi_{\star}\simeq 0.1$, where a quantity with a subscript $\star$ is evaluated at the loop production time $t_{\star}$. Consequently, the loop production function can be accurately approximated by
@@
where $\delta(x)$ denotes the Dirac $\delta$ function, which fixes the loop production time $t_{\star}$ through the relation $\alpha_{\mathrm{r}}\xi_{\star}={l_{\star}}/{t_{\star}}$. Substituting the expression for $\mathcal{P}_{\mathrm{r}}(l,t)$ into Eq. (22) with $a\propto t^{1/2}$ in the RD era, we derive an analytical expression for the loop number density [84, 80, 81],
@@
which agrees with the result from numerical simulations presented in Ref. [83].
On the other hand, numerical simulations for an MD era suggest that loop production occurs on a wider range of scales compared to that in an RD era and can be effectively described by a power-law distribution of $l^{-1.69}$ with a cutoff [83]. This motivate us to assume the loop production function as
@@
where the theta function gives a cutoff at $l=\alpha_{\mathrm{m}}L$ with $\alpha_{\mathrm{m}}$ satisfying $\alpha_{\mathrm{m}}\xi_{\star}\simeq 0.18$ [83]. Using $a\propto t^{2/3}$ in the MD era, we derive the loop number density
@@
which is also consistent with the simulation result in Ref. [83].
We now analyze the SGWB spectrum generated by CS loops within the standard $\Lambda\mathrm{CDM}$ cosmological model, utilizing the expressions (27) and (34) for the loop number density in the scaling regime. In the $\Lambda\mathrm{CDM}$ model, the MD era succeeds the RD era at the time of matter-radiation equality, $t_{\mathrm{eq}}=51.1\pm 0.8~{}\mathrm{kyr}$ [76]. Notably, CS loops formed during the RD era and surviving into the MD era also contribute to the loop number density in the MD era, as given by [83]
@@
In the following section, we will examine the effects of an additional EMD era inserted into the RD era, focusing on its impact on the scale factor, the loop number density, and the SGWB spectrum. This modification may violate the scaling behavior, necessitating the use of the more general expressions for the loop number density as in Eqs. (24) and (32).
@@ III.1 Early matter-dominated era
Figure 2: Sketch of the cosmic timeline featuring an EMD era. In chronological order, the universe evolves through the original RD era, the EMD era, the second RD era, and the final MD era.
@@
Now we discuss the evolution of the related energy densities. The radiation and matter energy densities, $\rho_{\mathrm{r}}$ and $\rho_{\mathrm{m}}$, scale with the scale factor $a$ as [85]
@@
The scale factor evolves as $a\propto t^{1/2}$ in an RD era and $a\propto t^{2/3}$ in an MD era. Starting from an initial temperature $T_{\mathrm{ini}}$ in the original RD era, the $Y$ particles constitute a matter component with energy density $\rho_{\mathrm{Y}}$, while the plasma of SM particles has energy density $\rho_{\mathrm{SM}}$. Consequently, the energy densities evolve as

### v3

@@ I Introduction
Modern cosmology originates from the study of the universe expansion and element formation [1]. Further observations have given rise to the standard $\Lambda\mathrm{CDM}$ cosmological model, which suggests that after inflation and reheating, the universe successively experiences a radiation-dominated (RD) era, a matter-dominated era (MD), and a dark-energy-dominated era. Main evidences involve the observations of the cosmic expansion [2], the big bang nucleosynthesis (BBN) [3, 4], the cosmic microwave background [5, 6, 7, 8], and the accelerating expansion in recent eras [9, 10].
Nevertheless, these observations hardly date back to pre-BBN eras. Therefore, it is essential to maintain an open perspective regarding the early history of the universe before BBN [11]. Various hypotheses beyond the standard cosmic history have been put forward, such as an early matter-dominated (EMD) era [12, 13, 14, 15, 16, 17, 18, 19], a kination-dominated era [20, 21, 22, 23, 24, 25], and an intermediate inflationary era [26, 27, 28, 29]. The verification of such novel hypotheses necessitates observational probes capable of accessing epochs prior to BBN, when the universe was opaque to photons and traditional electromagnetic detection methods are ineffective. Nonetheless, this challenge can be addressed through observations of gravitational waves (GWs), a new messenger first detected in 2015 [30]. Unlike electromagnetic radiation, GWs can propagate freely through space, preserving information from the early universe and allowing them to reach present day observatories.
@@
In this work, we explore how an SGWB spectrum from a preexisting CS network is modified by the EMD era that arises within the framework of the dark matter (DM) dilution mechanism [59, 60, 61, 62, 63, 64, 65, 66, 67, 68]. This mechanism addresses the DM overproduction problem by relying on significant entropy production from the decay of a long-lived particle, referred to as the “dilutor”. For the dilution mechanism to be effective, the dilutor must dominate the energy density of the universe for a finite period, thereby introducing an EMD era that interrupts the conventional RD era. To quantify the impact of the EMD era, we trace the evolution of the total energy density of the universe by solving the relevant Boltzmann equations. This allows us to derive the resulting SGWB spectrum, which reflects the influence of the EMD era on CS dynamics and GW emissions.
This paper is structured as follows. In Section II, we review the formation of CS loops and analyze their number densities during the RD and MD eras, which are essential for assessing the SGWB spectrum induced by CS loops. In Section III, we investigate an EMD era emerging from the DM dilution mechanism and examine its effects on the evolution of the CS network, the loop number densities, and the resulting SGWB spectrum. Finally, in Section IV, we summarize our findings.
@@ II Gravitational waves originating from cosmic strings
In a scalar field theory with a global or gauge $\mathrm{U}(1)$ symmetry, CSs could be formed after the spontaneous breaking of the $\mathrm{U}(1)$ symmetry in the early universe. They are one-dimensional topological defects concentrating the energy of the scalar field (and the gauge field for the gauged case) [69]. In the Nambu-Goto approximation, CSs are described as infinitely thin objects with tension $\mu$, which is the energy per unit length. The dimensionless quantity $G\mu$ is commonly used to represent the CS tension. Considering the dynamics of CSs, these combined factors establish CSs as promising sources of GWs [70, 35].
@@ II.1 Gravitational waves from cosmic string loops
Cosmic strings are generated randomly in the early universe, leading to the formation of the CS network. Long strings with super-horizon lengths intersect to form CS loops, whose relativistic oscillations can effectively emit GWs. Although small structures in long strings can also produce GWs, their contributions are generally negligible, compared to those generated by loops [71, 72, 37]. In practice, it is sufficient to consider only the stochastic GWs generated by CS loops.
@@
where $\Gamma$ is estimated to be approximately 50 [74], and $G$ is the Newtonian gravitational constant. The frequencies of GWs emitted by a CS loop of length $l$ are [70]
@@
where $P_{n}$ is the dimensionless emission power in units of $G\mu^{2}$ for a single mode $n$, which can be estimated by numerical simulations for GW emissions from the RD and MD eras [74].
@@
Note that this result corresponds to the GW emission time $t_{\mathrm{e}}$ but GW experiments only receive signals at the present time $t_{0}$. Hence, the effect of the cosmological redshift must be accounted for, and the present GW frequency is given by $f=a(t_{\mathrm{e}})f_{\mathrm{e}}$, while the present GW energy density is $\rho_{\mathrm{GW}}=\rho_{\mathrm{GW}}(t_{\mathrm{e}})a^{4}(t_{\mathrm{e}})$, where $a(t)$ is the scale factor normalized such that $a(t_{0})=1$. Integrating Eq. (5) over time, we arrive at
@@
is a function of the redshift $z$ introduced to account for the changes in relativistic degrees of freedom, where $g_{\star}(z)$ and $g_{\star s}(z)$ are the effective numbers of relativistic degrees of freedom for the energy and entropy densities, respectively. Considering the evolution of $g_{\star}(z)$ and $g_{\star s}(z)$, we can approximate $\mathcal{G}(z)$ as a piecewise function that changes at the epochs of electron-positron annihilation $(z\simeq 10^{9})$ and QCD phase transition $(z\simeq 2\times 10^{12})$[75]
@@
However, if an EMD era exists, the evolution of the Hubble rate $H(t)$ would be altered, necessitating modifications to the expressions above. This situation is discussed in detail in Section III.
@@ II.2 Number density of cosmic string loops
For calculating the SGWB spectrum induced by CS loops, estimating the loop number density distribution $n_{\mathrm{CS}}(l,t)$ is crucial. Two models can be used to evaluate $n_{\mathrm{CS}}(l,t)$ based on numerical simulations and the scaling nature of the CS network . The first is the BOS model, introduced by Blanco-Pillado, Olum, and Shlaer [77]. This model uses the horizon distance as the sole kinematic scale to describe the scaling behavior and extrapolates the loop production functions from numerical simulations to derive $n_{\mathrm{CS}}(l,t)$ in both RD and MD eras. The second is the LRS model, proposed by Lorenz, Ringeval, and Sakellariadou [78]. In this model, the number density distribution of produced loops per unit time is assumed to follow a power law in the scaling regime, and the gravitational backreaction effect is included to reduce loop production below a certain scale. Compared to the SGWB spectrum produced by the BOS model, the SGWB spectrum predicted by the LRS model has a significantly higher amplitude. Consequently, the constraint on the CS tension derived from the LIGO-Virgo O3 dataset is more stringent for the LRS model than for the BOS model [79]. Additionally, future space-borne interferometers will more easily detect the SGWB predicted by the LRS model [80].
Nevertheless, these two models rely on the scaling behavior of the CS network, which may be violated by the inclusion of an EMD era. To account for this nonscaling effect, we instead adopt the velocity-dependent one-scale (VOS) model [72, 81, 82], which describes the dynamical evolution of the CS network in a generic manner. This model has also been used to interpret the positive evidence of nHz SGWB reported by PTA experiments [83, 84, 85]. Below, we will demonstrate that the predictions of the VOS model align with those of the BOS model when the scaling regime is achieved and appropriate loop production functions are assumed. Furthermore, the VOS model can be directly applied to cosmic history with an EMD era, which is discussed in Subsection III.3.
In the VOS model, the dynamics of the CS network is characterized by two fundamental parameters, the correlation length $L$ and the root-mean-square (RMS) velocity $v$ of string segments. Thus, the energy density of long strings can be expressed as $\rho=\mu/L^{2}$, and its evolution in the universe is governed by [81]
@@
where the friction effect is neglected. The term $-2H\rho$ accounts for the dilution and stretching of the strings due to cosmic expansion, while $-2Hv^{2}\rho$ captures the energy loss caused by velocity redshift. The term $-\tilde{c}v\rho/L$ represents the energy removed through loop formation, where $\tilde{c}\simeq 0.23$ [86] is the loop chopping efficiency. Additionally, the evolution of the RMS velocity $v$ is determined by [81]
@@
Introducing the dimensionless quantity $\xi\equiv L/t$, which represents the correlation length $L$ normalized by the cosmic time $t$, the evolution equations above can be rewritten as
@@
By numerically solving these equations, it is found that the solutions rapidly converge to constant values of $\xi$ and $v$, independent of the initial conditions. This indicates that the CS network quickly evolves into a linear scaling regime characterized by $L\propto t$. By setting $\dot{\xi}=\dot{v}=0$, the scaling solutions for RD and MD eras can be derived as [87]
@@
where the Lorentz factor $\gamma_{v}\equiv(1-v^{2})^{-1/2}$ incorporates the energy loss caused by the redshifting of the loop velocity, and $\mathcal{F}$ is a coefficient that characterizes the fraction of slow, large loops responsible for the dominant contribution to GW emissions. The values of $\mathcal{F}$ are different for the RD and MD eras, denoted as $\mathcal{F}_{\mathrm{r}}$ and $\mathcal{F}_{\mathrm{m}}$, respectively. These values can be determined by a comparison with numerical simulations in the scaling regime [89, 77].
@@
quantifies the rate of increase in the number density of CS loops per unit length per unit time. Because of GW emission, a CS loop of length $l^{\prime}$ at time $t^{\prime}$ loses energy $\Gamma G\mu^{2}(t-t^{\prime})$ by time $t$ and shrinks to a length $l=l^{\prime}-\Gamma G\mu(t-t^{\prime})$. Therefore, the loop number density per unit loop length at time $t$ is given by
@@
with $l^{\prime}(t)=l+\Gamma G\mu(t-t^{\prime})$. $t_{\mathrm{ini}}$ denotes the initial time of loop production, and the scale factors account for the dilution caused by cosmic expansion. To determine $n_{\mathrm{CS}}(l,t)$ and calculate the energy density of the emitted GWs using Eq. (4), the loop production function $\mathcal{P}(l,t)$ is required.
During the RD era, numerical simulations [77] reveal that most large loops are generated with a length $l$ characterized by the fraction $\alpha_{\mathrm{r}}=l/L$ of the correlation length $L$. $\alpha_{\mathrm{r}}$ satisfies $\alpha_{\mathrm{r}}\xi_{\star}\simeq 0.1$, where the quantity with subscript $\star$ is evaluated at the loop production time $t_{\star}$. Consequently, the loop production function can be accurately approximated by
@@
where $\delta(x)$ denotes the Dirac $\delta$ function, which fixes the loop production time $t_{\star}$ through the relation $\alpha_{\mathrm{r}}\xi_{\star}={l_{\star}}/{t_{\star}}$. Substituting the expression for $\mathcal{P}_{\mathrm{r}}(l,t)$ into Eq. (22) with $a\propto t^{1/2}$ in the RD era, we derive an analytical expression for the loop number density [90, 87, 88]:
@@
which agrees with the result derived from numerical simulations in Ref. [77].
In contrast, numerical simulations for the MD era suggest that loop production occurs on a wider range of scales compared to that in the RD era and can be effectively described by a power-law distribution of $l^{-1.69}$ with a cutoff [77]. This motivate us to assume the loop production function as
@@
where the theta function has a cutoff at $l=\alpha_{\mathrm{m}}L$, with $\alpha_{\mathrm{m}}$ satisfying $\alpha_{\mathrm{m}}\xi_{\star}\simeq 0.18$ [77]. Using $a\propto t^{2/3}$ in the MD era, we derive the loop number density
@@
which is also consistent with the simulation results in Ref. [77].
We now analyze the SGWB spectrum generated by CS loops within the standard $\Lambda\mathrm{CDM}$ cosmological model, utilizing Eqs. (27) and (34) for the loop number density in the scaling regime. In the $\Lambda\mathrm{CDM}$ model, the MD era succeeds the RD era at the time of matter-radiation equality, $t_{\mathrm{eq}}=51.1\pm 0.8~\mathrm{kyr}$ [76]. Notably, CS loops formed during the RD era and surviving into the MD era also contribute to the loop number density in the MD era, as given by [77]
@@
In the following section, we will examine the effects of an additional EMD era inserted into the RD era, focusing on its impact on the scale factor, loop number density, and SGWB spectrum. This modification may violate the scaling behavior, necessitating the use of the more general expressions for the loop number density, as in Eqs. (24) and (32).
@@ III.1 Early matter-dominated era
Figure 2: Sketch of the cosmic timeline featuring an EMD era. In chronological order, the universe evolves through the original RD, EMD, second RD, and final MD eras.
@@
Now, we discuss the evolution of the related energy densities. The radiation and matter energy densities, $\rho_{\mathrm{r}}$ and $\rho_{\mathrm{m}}$, scale with the scale factor $a$ as [91]
@@
The scale factor evolves as $a\propto t^{1/2}$ in the RD era and $a\propto t^{2/3}$ in the MD era. Starting from an initial temperature $T_{\mathrm{ini}}$ in the original RD era, the $Y$ particles constitute a matter component with energy density $\rho_{\mathrm{Y}}$, while the plasma of SM particles has energy density $\rho_{\mathrm{SM}}$. Consequently, the energy densities evolve as